1) sandra spent 2/5 of an hour helping a customer extange an unwanted gift. then she spent 1/2 an hour giving a costumer a refund. what fraction of an hour did she spend helping both customer?
2)efrain the carpenter cut off 1/8 of a 56 inch board. how many inches did cut off?
3)two-fifths of the area of school parking lot can be used for a cultural festival. if the whole parking lot is 7800 square feet, how many square feet can be used for the festival?
4)sprockets R Us is selling its bicycle sprockets for 2/5 of a dollar. this price shows a decrease of 1/3 of the original price. what was the original price?
5)shad made 9 of the 25 free throws at basketball practice. what percent did he make?



Answer by edjones(8007)

2) 56 * 1/8 = 56/8 = 7"